### v1

# TwT: Thinking without Tokens by Habitual Reasoning Distillation with Multi-Teachers’ Guidance

Abstract
Large Language Models (LLMs) have made significant strides in problem-solving by incorporating reasoning processes. However, this enhanced reasoning capability results in an increased number of output tokens during inference, leading to higher computational costs. To address this challenge, we propose TwT (Thinking without Tokens), a method that reduces inference-time costs through habitual reasoning distillation with multi-teachers’ guidance, while maintaining high performance. Our approach introduces a Habitual Reasoning Distillation method, which internalizes explicit reasoning into the model’s habitual behavior through a Teacher-Guided compression strategy inspired by human cognition. Additionally, we propose Dual-Criteria Rejection Sampling (DCRS), a technique that generates a high-quality and diverse distillation dataset using multiple teacher models, making our method suitable for unsupervised scenarios. Experimental results demonstrate that TwT effectively reduces inference costs while preserving superior performance, achieving up to a 13.6% improvement in accuracy with fewer output tokens compared to other distillation methods, offering a highly practical solution for efficient LLM deployment.
TwT: Thinking without Tokens by Habitual Reasoning Distillation
with Multi-Teachers’ Guidance
Jingxian Xu1thanks:   Work during internship at Microsoft., Mengyu Zhou4thanks:   Corresponding author (mezho@microsoft.com)., Weichang Liu211footnotemark: 1, Hanbing Liu311footnotemark: 1, Shi Han4, Dongmei Zhang4 1Nankai University 2BeijingJiaoTong University 3Tsinghua University 4Microsoft Research
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable improvements in problem-solving by incorporating reasoning process Brown et al. (2020); Wei et al. (2022); Chowdhery et al. (2023); Yao et al. (2024). It enhances the reasoning capability of LLMs by breaking down complex tasks into intermediate steps, leading to better performance. However, reasoning capability comes at a significant cost: the reasoning process substantially increases the number of output tokens during inference, resulting in higher inference-time computational costs Snell et al. (2024); Wu et al. (2024). In practical deployments, computational resources and budgets are often constrained, making the reduction of inference-time computational costs a pressing issue that requires effective solutions Zheng et al. (2022); Chung et al. (2024).
Figure 1: Overview of the Proposed Method. The upper part of the figure illustrates the background problem: while generating more reasoning steps improves the performance of LLMs, it also leads to significantly higher computational costs. To mitigate this, we propose targeted strategies, shown in the lower part of the figure. Our approach reduces the cost per token by distilling knowledge from large models into smaller ones and minimizes the total number of tokens by gradually shrinking intermediate reasoning paths.
Recent research has primarily explored two key approaches to addressing this issue: reducing the cost per token and reducing the number of reasoning tokens as illustrated in Figure 1 Wang et al. (2024a); Hsieh et al. (2023); Wang (2024). A common strategy for reducing the cost per token is to replace large models with smaller, more efficient ones. However, directly using smaller models often results in a significant drop in performance, particularly on complex reasoning tasks. To mitigate this problem, knowledge distillation Hinton (2015); Park et al. (2019) has emerged as an effective solution, enabling student models to mimic the performance of larger teacher models Tang et al. (2019); Hsieh et al. (2023). However, traditional knowledge distillation typically relies on expensive human-labeled data, limiting its practical applicability. Additionally, most distillation methods employ only a single teacher model, restricting the diversity of knowledge that could be leveraged from multiple teacher models. To overcome these limitations, we propose Dual-Criteria Rejection Sampling (DCRS), a method that first utilizes a multi-teacher strategy to generate pseudo-labels and then applies a two-stage selection process-Quality Selection and Diversity Selection-to construct a high-quality and diverse distillation dataset. This approach not only enhances the efficiency of knowledge transfer but also enables our model to adapt effectively to unsupervised settings.
For the other issue, the number of reasoning tokens can be reduced by shrinking intermediate reasoning paths Wang et al. (2024a); Deng et al. (2024). While this effectively lowers token usage, it often comes at the cost of degraded model performance. Therefore, it is essential to develop a method that balances computational efficiency and model accuracy. Consider a real-world learning scenario where a teacher possesses a deep and comprehensive understanding of a concept, while a student may struggle to grasp the material fully. To bridge this gap, the teacher distills knowledge, extracting only the most essential and concise information to help the student learn more effectively. Over time, the student internalizes this reasoning process, enabling them to generate answers instantly upon encountering a question, without requiring explicit intermediate reasoning steps.
Inspired by this human cognitive process, we propose Habitual Reasoning Distillation (HaRD), a method that internalizes explicit reasoning into the model’s habitual behavior through a multi-stage distillation process, thereby reducing the need for explicit reasoning during inference. HaRD follows a three-stage distillation strategy: (a) Full Reasoning Distillation, where the student learns reasoning patterns from complete reasoning paths generated by teacher models; (b) Reasoning-Compressed Distillation, where the reasoning process is progressively compressed, with teachers refining their outputs based on the student’s responses to create reasoning paths aligned with the student’s capabilities; and (c) Reasoning-Free Distillation, where the student is trained without explicit reasoning steps, relying only on final labels, allowing it to generate high-quality answers directly. This process shifts the computational burden from inference to training, enabling both high performance and low inference cost.
In this work, we propose TwT (Thinking without Tokens), a method that achieves an optimal balance between inference-time computational cost and performance. TwT follows a two-step process. First, DCRS utilizes multi-teacher LLMs to generate pseudo-labels, enabling the model to adapt to unsupervised settings. Then, HaRD applies a multi-stage distillation approach to progressively internalize explicit reasoning abilities into the student model as inherent capabilities. Our key contributions are summarized as follows:
  • Novel Distillation Framework: We propose TwT, a novel framework that aims to reduce inference-time computational cost through habitual reasoning distillation with multi-teachers’ guidance while preserving high performance.
  • Unsupervised Sampling Strategy: We propose Dual-Criteria Rejection Sampling (DCRS), a method that selects high-quality and diverse distillation data generated by multi-teacher LLMs, enabling adaptation to unsupervised settings.
  • Efficient Reasoning Distillation: We design a Habitual Reasoning Distillation (HaRD) method that refines reasoning patterns through a teacher-guided compression strategy, ensuring better alignment with the student model’s capabilities and ultimately integrates explicit reasoning into the model’s inherent behavior.
  • Comprehensive Empirical Validation: Experimental results demonstrate that our approach outperforms existing distillation techniques, achieving up to a 13.6% improvement in performance while generating fewer tokens.
Figure 2: Method Framework. Our proposed TwT (Thinking without Tokens) framework consists of two stages: Dual-Criteria Rejection Sampling (DCRS) and Habitual Reasoning Distillation (HaRD). In the first stage, DCRS selects a high-quality and diverse reasoning distillation dataset generated by multiple teacher LLMs (e.g., T1, T2, T3). In the second stage, HaRD progressively internalizes reasoning ability into the student model through a three-stage distillation process.
2 Related Work
2.1 Knowledge Distillation for LLMs
Knowledge distillation (KD) Hinton (2015); Beyer et al. (2022); West et al. (2021) transfers capabilities from large LLMs to smaller ones using hard or soft labels from a teacher model. However, by relying solely on final outputs, standard KD provides limited information. Reasoning distillation Hsieh et al. (2023) addresses this by training students to understand both the final answer and the underlying reasoning. Recent work Chen et al. (2023); Liu et al. (2023) explores generating multiple rationales per query to enhance robustness. However, using a single teacher can introduce bias and limit diversity. Multi-teacher strategies Tian et al. (2024); Zhang et al. (2024) address this by aggregating diverse reasoning paths, enriching distillation data, and improving generalization. These methods rely on labeled datasets to select high-quality distilled data, which can be difficult to obtain. Our work introduces a multi-teacher approach to incorporate diverse reasoning data and proposes a Dual-Criteria Rejection Sampling strategy to obtain a high-quality and diverse distillation dataset from unlabeled data.
2.2 Reasoning and Inference-time Scaling
Recent work enhances output diversity by improving reasoning via structural methods (e.g., code parsing, problem decomposition Gao et al. (2023); Zhou et al. (2022)) and by generating multiple reasoning paths through techniques like majority voting and reinforcement learning Wei et al. (2022); Yao et al. (2023); Cao (2024); Wang et al. (2022); Fu et al. (2022); Huang et al. (2023); Trung et al. (2024); Wang et al. (2024b). However, these supervised approaches rely on a single model and labeled data, limiting inherent diversity. Additionally, studies such as Snell et al. (2024); Wu et al. (2024) show that adaptive inference-time strategies can significantly reduce compute costs, but model performance will be reduced. To address these issues, we propose a multi-teacher strategy, an unsupervised approach for generating diverse, high-quality samples with habitual reasoning distillation for efficient inference with explicit reasoning.
3 Method
In this section, we provide a detailed explanation of the implementation of our TwT as illustrated in Fig. 2. First, we propose a Dual-Criteria Rejection Sampling (DCRS) strategy to obtain high-quality and diverse distillation samples (Section 3.1). Then, we design a Habitual Reasoning Distillation (HaRD) strategy that progressively internalizes the reasoning ability at each distillation stage, allowing the reasoning capabilities of the teacher models to be gradually internalized into the student model (Section 3.2).
Figure 3: Dual-Criteria Rejection Sampling Architecture. Our proposed DCRS method comprises two stages: Quality Selection and Diversity Selection. The first stage filters samples using confidence scores, while the second stage enhances diversity by selecting samples based on similarity. This approach ensures a high-quality and diverse distillation dataset, enabling our method to effectively adapt to unsupervised scenarios.
3.1 Dual-Criteria Rejection Sampling
To provide the student model with high-quality and diverse reasoning paths, we propose a novel paradigm termed Dual-Criteria Rejection Sampling (DCRS), which extends traditional rejection sampling Gilks and Wild (1992) by integrating two key selection metrics: confidence scores and similarity measures. Leveraging a multi-teacher strategy, we first prompt teacher LLMs to generate an initial pool of pseudo-labels. DCRS performs sample selection in two sequential steps: quality selection and diversity selection.
3.1.1 Quality Selection
As shown in Figure 3, we take use of CoT prompting Wei et al. (2022) to generate and extract reasoning patterns from multi-teacher LLMs. Given an unlabeled dataset $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$, where each $x_{i}$ is a query, we first design a prompt template $p$ to clarify the task solution method. The prompt instructs the LLM to produce the output $O_{i}$ in the form of a triplet $(r_{i},y_{i},c_{i})$, where $y_{i}$ is the predicted label for task $x_{i}$, $r_{i}$ is the rationale provided by the LLM, and $c_{i}$ is the confidence score for the reasoning and predicted label, represented as a decimal in the range $[0,1]$. To ensure that the confidence score is more reliable, we compute it using a weighted combination of multiple performance factors:
$c_{i}=\sum_{j=1}^{n}w_{j}\cdot m_{j}$ (1)
where $w_{j}$ denotes the weight assigned to the $j$-th factor and $m_{j}$ represents the corresponding factor value. For example, in a code generation task, we let the large model assign a weighted score based on factors such as the reasoning process, code readability, and robustness, yielding the final confidence score. For $k$ teacher models, we can obtain a set of outputs $O_{i}=\{O_{i1},O_{i2},\dots,O_{ik}\}$ for each query $x_{i}$. Then, we set a confidence score threshold $s$. For each output $O_{i}$, if its confidence score $c_{i}\geq s$, the output sample is considered high-quality and retained; otherwise, it is discarded. Subsequently, diversity selection is performed on the retained sample set =$\mathcal{H}=\{(x_{i},r_{i},y_{i},c_{i})\}_{i=1}^{P}$, where $P$ is the total number of retained samples.
3.1.2 Diversity Selection
For the high-quality sample set $\mathcal{H}$, if the outputs $O_{i}=\{O_{i1},O_{i2},\dots,O_{ik}\}$ for the same query $x_{i}$ come from three or more different teacher models (i.e., $k\geq 3$), we calculate the semantic similarity between the rationales provided by these teacher models. Specifically, we assume that $r_{ij}$ denotes the rationale generated by the $j$-th teacher for query $x_{i}$. Each rationale $r_{ij}$ is then mapped into a fixed-dimensional embedding $E(r_{ij})$ using a pre-trained sentence embedding model. The cosine similarity between any two rationales $r_{ip}$ and $r_{iq}$ is computed as:
$\text{sim}(r_{ip},r_{iq})=\frac{E(r_{ip})\cdot E(r_{iq})}{\|E(r_{ip})\|\,\|E(r_{iq})\|}$ (2)
where $r_{ip}$ and $r_{iq}$ are rationales provided by two distinct teacher models for the same question, with $1\leq p<q\leq k$. We calculate the cosine similarity for all unique pairs $(r_{ip},r_{iq})$. To maximize diversity, we select the pair of rationales that yields the lowest similarity score. In our implementation, we directly select the pair of rationales that exhibits the minimum cosine similarity. For a given query $x_{i}$, we define:
$(p^{*},q^{*})=\arg\min_{1\leq p<q\leq k}\text{sim}\bigl{(}r_{ip},r_{iq}\bigr{)})$ (3)
where $(p^{*},q^{*})$ are the indices corresponding to the pair of rationales with the smallest similarity, thereby ensuring that the final distillation dataset $\mathcal{G}=\{(x_{i},r_{i},y_{i})\}_{i=1}^{Q}$ is both high-quality and diverse, where $P$ is the total number of final distillation samples.
3.2 Habitual Reasoning Distillation
In this section, we propose a novel distillation strategy to address the trade-off between inference-time computational cost and model performance. Our approach decomposes the distillation process into three sequential stages. In Stage-1, the student model is distilled using data with full reasoning, allowing it to fully absorb the teacher’s comprehensive thinking process. During the Stage-2, the model learns from data with compressed reasoning, encouraging it to derive solutions using more concise and efficient explanations. Finally, in Stage-3, distilling is conducted on data without any explicit reasoning, enabling the student to learn an end-to-end mapping from query to answer. Notably, in the second stage, we propose a Teacher-Guided Compression method that tailors the complexity of the reasoning information to the student model’s capacity.
Stage-1: Full Reasoning Distillation.
In this stage, the student model is trained to learn the complete reasoning paths under the supervision of the teacher models’ full reasoning. The goal is to help the weak student model understand the logical steps involved in the task and build a solid foundation for further distillation.
Method MBPP CQA MetaMath
Pass@1 Token Accuracy Token Accuracy Token
GPT-4 77.68% 493 83.27% 312 86.31% 512
GPT-4o-mini 77.14% 451 81.76% 301 87.03% 529
Mistral-large 73.83% 365 80.15% 288 86.33% 463
Mistral-7B-v0.3 Vanilla Student 42.90% 209 60.19% 171 30.12% 255
Standard KD 52.67% 53 63.63% 5 38.19% 7
Distilling 53.39% 170 64.47% 129 42.20% 397
Tinyllm 54.45% 175 67.39% 225 41.67% 226
TwT 57.11% ( $\uparrow 4.88\%$ ) 48 76.16% ( $\uparrow 13.01\%$ ) 6 47.94% ( $\uparrow 13.60\%$ ) 7
Phi-3.5mini Vanilla Student 54.71% 199 64.15% 150 73.36% 279
Standard KD 61.71% 98 64.48% 6 78.16% 10
Distilling 62.03% 202 65.74% 148 78.19% 370
Tinyllm 64.44% 192 70.30% 144 78.23% 240
TwT 67.93% ( $\uparrow 5.42\%$ ) 105 76.42% ( $\uparrow 8.70\%$ ) 10 83.58% ( $\uparrow 6.84\%$ ) 15
Table 1: Quantitative results for baseline models. The top three rows show the inference results of our teacher models, while “Distilling” is an abbreviation for “Distilling Step-by-Step.” The best and the second best results are highlighted in bold and underlined respectively. The improvements of TwT over the second best results are shown in green with an upward arrow.
The teachers’ reasoning ability can be transferred by fine-tuning the student model using the full reasoning $\mathcal{G}$ derived from high-quality and diversity sample selection. More specifically, the process of learning full reasoning paths through fine-tuning is defined as follows:
$\mathcal{L}_{\mathrm{1}}=\mathbb{E}_{\mathcal{G}}[\log\mathrm{P}_{f}([x;r;y])]$ (4)
where $f$ indicates the student model, $\mathbb{E}_{\mathcal{G}}$ is the expectation over the distillation dataset $\mathcal{G}$, and $\mathrm{P}_{f}([x;r;\hat{y}])$ is the probability assigned by the student model f to the joint input $[x;r;\hat{y}]$.
Stage-2: Reasoning-Compressed Distillation.
In this stage, we progressively simplify the reasoning paths of the teacher model, generating more concise one by compressing the original reasoning paths. We observed that for the same problem, the outputs of student models are often shorter and feature more concise reasoning steps compared to those of teacher models. Therefore, we adopt a Teacher-Guided Compression approach that ensures the reasoning paths provided by the teacher models are better aligned with the learning characteristics of the student model, thereby enhancing overall distillation performance.
Specifically, for a given query $x_{i}$, the teacher model generates the original reasoning $r_{i}$, while the student model produces the reasoning $r^{s1}_{i}$. We design a prompt $p^{\prime}$ to guide the teacher model in refining its original reasoning $r_{i}$ based on the characteristics of the student model’s output (e.g., output length, complexity of understanding the problem). This process can be represented as $(p^{\prime},r_{i},r^{s1}_{i})$ $\rightarrow r^{T}_{i}$. Subsequently, we replace the original reasoning $r_{i}$ in the dataset $\mathcal{G}=\{(x_{i},r_{i},y_{i})\}_{i=1}^{N}$ with the refined reasoning $r^{T}_{i}$, resulting in the second-stage distillation dataset $\mathcal{G^{\prime}}=\{(x_{i},r^{T}_{i},y_{i})\}_{i=1}^{N}$. The second stage fine-tuning can be defined as:
$\mathcal{L}_{\mathrm{2}}=\mathbb{E}_{\mathcal{G^{\prime}}}[\log\mathrm{P}_{f}([x;r^{T};y])]$ (5)
Stage-3: Reasoning-Free Distillation.
Finally, we completely remove the reasoning paths and only retain the final answer as the supervision signal. The student model is trained to directly output the correct answer without relying on any reasoning chain. The goal of this stage is to enable the student model to form a “habitual” ability, allowing it to efficiently complete tasks without the need for complex reasoning. After removing $r$, the new dataset $\mathcal{G^{\prime\prime}}=\{(x_{i},y_{i})\}_{i=1}^{N}$ only contains the original query $x_{i}$ and the label $y_{i}$ predicted by the teacher model. The third stage fine-tuning can be defined as:
$\mathcal{L}_{\mathrm{3}}=\mathbb{E}_{\mathcal{G^{\prime\prime}}}[\log\mathrm{P}_{f}([x;y])]$ (6)
In summary, Stage-1 and Stage-2 use full resoning distillation and reasoning-compressed distillation, respectively, to gradually familiarize the student model with the complete reasoning chain and establish a systematic reasoning pattern. This enables the model to master the key problem-solving steps at both conceptual and operational levels, ultimately internalizing complex reasoning abilities and forming a stable reasoning mechanism. In Stage-3, we perform end-to-end training to further deepen the model’s understanding of the relationship between problems and their answers. Since the explicit reasoning chain has been embedded as an inherent capability in the previous stages, Stage-3 no longer relies on explicit reasoning paths but instead directly reinforces answer generation. This progressive strategy allows the student model to output correct answers without explicit reasoning, thereby reducing inference-time computational cost while enhancing overall performance.
4 Experiment
Method MBPP CQA MetaMath
Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini
Pass@1 Token Pass@1 Token Accuracy Token Accuracy Token Accuracy Token Accuracy Token
TwT-stage1 54.83% 291 65.32% 310 73.31% 141 72.99% 138 46.48% 295 80.05% 313
TwT-stage2 56.48% 154 66.42% 184 74.89% 84 75.39% 73 47.02% 169 81.33% 196
TwT-stage3 57.11% 48 67.93% 105 76.16% 8 76.42% 10 47.94% 12 83.58% 15
Table 2: Quantitative results for three distillation stages. Accuracy and the number of output tokens were used to evaluate the model performance. The best results were highlighted in bold.
Method MBPP CQA MetaMath
Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini
Pass@1 Token Pass@1 Token Accuracy Token Accuracy Token Accuracy Token Accuracy Token
TwT-stage1 54.83% 291 65.32% 310 73.31% 141 72.99% 138 46.48% 295 80.05% 313
TwT-stage2 56.48% 154 66.42% 184 74.89% 84 75.39% 73 47.02% 169 81.33% 196
TwT-stage3 56.94% 133 67.18% 161 75.44% 55 75.85% 64 47.41% 144 81.99% 172
TwT-stage4 57.49% 42 68.21% 100 76.29% 8 76.88% 9 48.37% 12 83.61% 13
Table 3: Quantitative results for four distillation stages. Accuracy and the number of output tokens were used to evaluate the model performance. The best results were highlighted in bold.
4.1 Experiment Setup
Datasets. We evaluate our TwT on 3 benchmark datasets for 3 different NLP tasks: MBPP Austin et al. (2021) for NL to python code generalization; CommonsenseQA (CQA) Talmor et al. (2018) for commonsense question answering; MetaMathQA (MetaMath) Yu et al. (2023) for mathematical reasoning, which is augmented from the training sets of GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021).
Models. We utilize GPT-4, GPT-4omini and Mistral-Large as our teacher models, which are accessed through OpenAI’s API and MistralAI’s API. For the student models, we use Mistral-7B-v0.3 and Phi-3.5mini. For the pre-trained sentence embedding model, we leverage all-mpnet-base-v2.
Baselines. For our baselines, we evaluate three types of methods: teacher model’s performance, vanilla student model’s performance, and knowledge distillation based methods that containing Standard-KD Hinton (2015), a general distillation method that fine-tunes the student model using the teacher model’s generated labels as ground-truth; Distilling-Step-by-Step Hsieh et al. (2023), which leverages LLM-generated rationales as additional supervision to train smaller models; TinyLLM Tian et al. (2024), a paradigm that distills diverse reasoning paths from multiple teacher LLMs into a student model.
Implementation Details. In our experimental setup, we employed training with LoRA fine-tuning, setting the LoRA rank to 8, a learning rate of 1e-5, a batch size of 8, 4 training epochs, and a context window of 4096 tokens. During inference, we used a temperature of 0, max tokens set to 2048, and a top-p value of 0.95. For the sampling process, we selected a scoring threshold of $s=0.95$. All experiments were conducted on four NVIDIA A100 Tensor Core GPUs, enabling large-scale training and efficient computation. The prompts for specific methods and a case study are provided in the appendix.
4.2 Baseline Comparison
Across three specific-tasks, TwT consistently outperforms other distillation methods as shown in Table 1. Compared with the best performing baseline, TwT achieves an improvement of up to 13.60% compared to the best-performing baseline, while reducing the token number by 98.2% (token numbers from 397 to 7 on MetaMath dataset), substantially lowering the inference cost. Typically, reducing inference tokens leads to a decline in performance; however, TwT overcomes this trade off by maintaining or even enhancing model performance while dramatically reducing token usage, thus achieving both high performance and low inference-time computational cost simultaneously. In addition, TwT effectively bridges the gap between the student and teacher models, significantly narrowing the performance disparity observed in vanilla student models.
Samping Method MBPP CQA MetaMath
Log Probability-Based 74.83% 75.29% 72.19%
Hard Rejection Sampling-Based 72.29% 73.11% 70.26%
Confidence Score-Based 83.49% 84.90% 81.14%
Table 4: Quantitative results for sampling methods. Accuracy was used to evaluate the model performance. The best results were highlighted in bold.
Compression Method MBPP CQA MetaMath
Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini Mistral-7B-v0.3 Phi-3.5mini
Fixed-Length Compression 55.89% 64.94% 73.29% 72.92% 43.18% 80.62%
Compressor 56.01% 65.10% 73.61% 73.69% 45.22% 80.95%
Teacher-guided Compression 56.48% 66.42% 74.89% 75.39% 47.02% 81.33%
Table 5: Quantitative Results for Compression Methods. Accuracy was used to evaluate the model performance. The best results were highlighted in bold.
Methods MBPP CQA MetaMath
Accuracy Token Accuracy Token Accuracy Token
w/o Multi-Teacher Strategy (GPT-4) 55.38% 54 74.64% 7 46.82% 15
w/o DCRS 55.42% 48 74.79% 9 47.11% 14
w/o Compression Distillation Stage 54.49% 93 72.19% 21 46.38% 48
TwT 57.11% 48 76.16% 8 47.94% 12
Table 6: Ablation Study for Model Components. Accuracy and the number of output tokens were used to evaluate the model performance. The best results were highlighted in bold.
4.3 Distillation Stage Analysis
In the Habitual Reasoning Distillation phase, we separately evaluated the student model’s performance at each fine-tuning stage and tracked the number of output tokens during inference, as shown in Table 2. The results indicate that TwT steadily improves with each stage, while the inference token numbers gradually decrease. By leveraging our distillation strategy, the model successfully internalizes the reasoning process as part of its own capabilities.
Furthermore, we expanded our three-stage procedure by extending it to four stages, adding an additional step after the second stage where the teacher model further compresses the reasoning process based on the student’s output. As shown in Table 3, this extension yields a slight improvement but does not significantly surpass the three-stage method. This suggests that our original three-stage process is already effective in achieving the desired performance.
4.4 Sampling and Compression Analysis
We evaluated the effectiveness of our DCRS method’s confidence score selection for distillation data and our Teacher-Guided Compression approach for enhancing student model’s performance. For sampling, we compared Confidence Score-Based sampling with two alternatives—Log Probability-Based selection and Hard Rejection Sampling, which categorizes LLM outputs into Good, Medium, and Poor—across three datasets. As shown in Table 4, our method consistently outperforms the alternatives with nearly a 10% gain in accuracy improvement. This highlights the effectiveness of our strategy in identifying high-quality examples for distillation. For compression, we compared our Teacher-Guided Compression with Fixed-Length Compression and a Compressor-based approach. As detailed in Table 5, our method better aligns the teacher’s outputs with the student model’s limited capacity, enhancing both training efficiency and overall performance.
4.5 Ablation Study
We further analyze the contribution of each component to TwT’s performance through ablation studies. Specifically, w/o Multi-Teacher Strategy evaluates the effect of using a single teacher model, w/o DCRS evaluates performance without filtering the distillation data, and w/o Compression Distillation Stage analyzes the impact of directly removing the reasoning step as shown in Table 6.
When comparing TwT with the multi-teacher strategy against the single-teacher approach, we observe an average improvement of 1.4% in the distillation performance. This result highlights the effectiveness of the multi-teacher strategy, as it provides a more diverse range of reasoning paths for the student model, thereby facilitating its learning process. Comparing TwT’s DCRS strategy to directly organizing the raw pseudo-labeled data genereated by teachers, TwT achieves an average improvement of 1.5%, underscoring the importance of performing high-quality and diverse data sampling before distillation. This step mitigates the influence of low-quality data and boosts overall distillation performance. Finally, when comparing TwT’s multi-stage distillation approach with the strategy of simply removing the reasoning process (i.e., without any intermediate compression), TwT achieves an approximate 3.5% accuracy gain, while the uncompressed approach not only underperforms in accuracy but also increases the number of output tokens. This finding confirms the necessity of gradually internalizing the reasoning capability for optimal performance.
5 Conclusion and Future Work
We introduced TwT, a novel distillation framework that internalizes reasoning abilities into a student model under multi-teachers’ guidance. It incorporates a Dual-Criteria Rejection Sampling stage to obtain a high-quality and diverse distillation dataset and a Habitual Reasoning Distillation strategy to gradually integrate reasoning capabilities into the student model. TwT achieves high performance with low inference cost without relying on labeled data or an explicit reasoning process. In future work, we will continue to explore whether further subdividing the distillation stages can enhance our framework
Limitations
Although our method has achieved excellent results, there are still some minor flaws here. One limitation of our approach is that it currently only works effectively on specific tasks and is not applicable to datasets containing mixed tasks. Additionally, the Dual-Criteria Rejection Sampling process could consist of noise. The impact of this potential noise on performance is still undetermined. A potential future direction is to investigate implicit natural language reasoning by utilizing more advanced training strategies. While current tasks are primarily focused on explicit reasoning, incorporating implicit reasoning mechanisms could improve the model’s robustness and its ability to generalize across different tasks.
Appendix A Prompt
A.1 Prompt for Teachers to Generate Pseudo-Labels on the MBPP, CQA and MetaMath Dataset
Figure 4: Prompt for Teachers to Generate Pseudo-Labels on the MBPP Dataset.
Appendix B Case Study on MBPP Dataset
B.1 Query
A specific query example of the MBPP dataset is shown below:
Figure 5: An example query of the MBPP dataset, showing the input programming task and its corresponding test case.
B.2 Teacher’s Original Response (HaRD Stage-1’s dataset)
After Dual-Criteria rejection sampling, one of the selected teacher responses (with full reasoning process) is:
Figure 6: The original structured response generated by the teacher model based on the query. The italicized part represents the full reasoning process, which will later be refined based on the student model’s response.
B.3 Student’s Inference for The Same Query
After performing SFT on the student model using the full reasoning dataset from the first stage, an initial student model with reasoning ability is obtained. Let this stage’s student model perform inference on the distillation dataset, resulting in the student’s response, as shown in Fig.7
Figure 7: The inference results of the student model after the first stage of SFT on the distill dataset.
B.4 Teacher’s Refinement Prompt
After the student model performs inference, the reasoning processes of both the student and teacher models are extracted for the same query. The teacher model modifies its original reasoning process based on the student model’s response preferences. The prompt is as follows:
Figure 8: Prompt for teacher’s refinement.
B.5 Compressed Reasoning Data After Refinement (HaRD Stage-2’s dataset)
By having the teacher model refine its original response based on the student model’s answer to the same question, the teacher’s response can better align with the student’s learning characteristics. This can be compared to the human learning process: the teacher delivers the initial lesson, then, after receiving feedback from the student, modifies the lesson plan based on the feedback to improve subsequent teaching. The teacher’s revised response is shown in the Fig.9. Since the refined reasoning is intuitively much shorter than the original reasoning, we refer to it as the “reasoning-compressed dataset”, which is used for the second stage of distillation in HaRD.
Figure 9: Reasoning-compressed distillation stage data.
B.6 Reasoning-free Distillation Data
After the second stage of SFT, the student model gradually internalizes the reasoning process into its own capability. In the third stage, we aim to enable the student model to learn query-answer end-to-end, strengthening its memory of the answer. Through this repetitive training, the student model will eventually be able to provide accurate answers without outputting reasoning. The format of the distillation data in the third stage is: prompt + query + answer (python code), meaning that the reasoning process is entirely removed. This dataset is used to continue SFT on the student model from the second stage, resulting in the final trained student model, completing the entire HaRD training process.
Appendix C Example Appendix
This is an appendix.